# Unified Domain Adaptive Semantic Segmentation

### v2

@@ Abstract
Unsupervised Domain Adaptive Semantic Segmentation (UDA-SS) aims to transfer the supervision from a labeled source domain to an unlabeled target domain. The majority of existing UDA-SS works typically consider images whilst recent attempts have extended further to tackle videos by modeling the temporal dimension. Although the two lines of research share the major challenges – overcoming the underlying domain distribution shift, their studies are largely independent. This causes several issues: (1) The insights gained from each line of research remain fragmented, leading to a lack of a holistic understanding of the problem and potential solutions. (2) Preventing the unification of methods, techniques, and best practices across the two domains, resulting in redundant efforts and missed opportunities for cross-pollination of ideas. (3) Without a unified approach, the knowledge and advancements made in one domain (images or videos) may not be effectively transferred to the other, leading to suboptimal performance and slower progress. Under this observation, we advocate unifying the study of UDA-SS across video and image scenarios, enabling a more comprehensive understanding, synergistic advancements, and efficient knowledge sharing. To that end, we explore the unified UDA-SS from a general data augmentation perspective, serving as a unifying conceptual framework, enabling improved generalization, and potential for cross-pollination of ideas, ultimately contributing to the overall progress and practical impact of this field of research. Specifically, we propose a Quad-directional Mixup (QuadMix) method, characterized by tackling distinct point attributes and feature inconsistencies through four-directional paths for intra- and inter-domain mixing in a feature space. To deal with temporal shifts with videos, we incorporate optical flow-guided feature aggregation across spatial and temporal dimensions for fine-grained domain alignment. Extensive experiments show that our method outperforms the state-of-the-art works by large margins on four challenging UDA-SS benchmarks. Our source code and models will be released at https://github.com/ZHE-SAPI/UDASS.
@@
Unified domain adaptation, semantic segmentation, QuadMix, flow-guided spatio-temporal aggregation.I Introduction
Semantic segmentation attempts to make precise dense predictions relevant to assigning a unique label to every single pixel within images or video frames, vital for a broad range of critical applications, such as autonomous driving [1, 2], medical imaging [3, 4], and so forth. However, when adapting to unseen and unlabeled video or image domains, the model performance always drops largely due to significant domain gaps at spatial [5] or temporal [6] levels, which limits their usability and deployability. Moreover, manual pixel-level annotation is labor-intensive [7], while labels of synthetic data rendered by game engines [8, 9, 10] are freely available. This has led to growing interest in Unsupervised Domain Adaptive Semantic Segmentation (UDA-SS), which transfers knowledge from a labeled source domain (synthetic) to an unlabeled target domain (real). However, despite the shared challenges in video (UDA-VSS) and image (UDA-ISS) scenarios, research in these areas is largely independent, leading to a lack of holistic understanding of the problem and potential solutions.
At present, there is not so substantial research on UDA-VSS. The community pays more attention to UDA on image scenarios, plenty of works [11, 12, 13, 14, 15, 16, 17, 18, 19, 5, 20, 21, 22, 23] address syn2real adaptation by joint learning. The methods for UDA-ISS can be roughly divided into two lines. The first line is based on domain alignment to reduce feature distribution discrepancies explicitly or implicitly. In practice, intermediate domain bridging [19, 21, 22, 23, 24], adversarial training [25, 26, 27], and statistical measure minimizing [28, 29, 30] have been proposed to diminish the domain gap. Despite the fact that they can pull the target domain distribution to the source domain, it does not guarantee the domain gap within each category-aware feature space can be solved, because pixels across domains could share similar semantics but feature quite distinct visual characteristics. Another line is self-supervised learning, adpoting confidence filtering [5, 31], label denoising [32, 33] or regularization [14, 34]. However, obtaining pseudo-labels is not easy when large gaps exist between distinct domains and always only denoised or high-confidence labeled samples are involved in self-supervised learning.
Naturally, existing UDA-ISS methods are not directly suitable for videos, owing to domain gaps not just at spatial but also temporal levels [35], and the object motion is complex [36]. Recently, several UDA-VSS methods based on self-supervised video learning or domain alignment strategies have been proposed [37, 36, 38]. In practice, they employ fine-designed auxiliary clues involving optical flow or geometric motion to constrain pixel-class consistency among adjacent frames. Moreover, techniques like one-way mixed intermediate video domain [39] stage-wisely or adversarial training [35, 6] have emerged to align feature distribution across domains. Nevertheless, similar to UDA-ISS works, the persistent domain gap in spatio-temporal feature embeddings hindered the knowledge transfer and reliable pseudo-label generation.
However, fragmented insights from independent scenarios often lead to redundant efforts and missed opportunities for cross-pollination, resulting in slower progress and suboptimal performance. From a unified view, current state-of-the-art methods achieve impressive results by addressing domain gaps through meticulously designed intermediates for image [19, 21, 22, 23, 24] and video [39] scenarios. However, we find that the compact nature of features within each domain (intra-domain) causes source and target domains to be represented as two distinct points [40], resulting in the intermediates with poorly generalized distributions for hard gap bridging. As shown in Fig. 1(a)-(d), we present a comparison of existing methods, distinguishing source features in green and target features in purple. Inspired by [41, 42], strategies like one-way mixing (e.g.,[19, 24, 23, 39]) or two-way mixing (e.g.,[21, 22]) extensively explore the integration of source pixel-level patches into the target domain, or vice versa, aiming to bridge the domain gap explicitly. Nonetheless, as supported by the t-SNE analysis [43] in Fig. 7(a)(c)(d), the effect of existing one-/two-way mixed intermediates across domains (inter-domain) that derived from such distinct domains is fragmentary and insufficient, which continues to hinder knowledge transfer to the target domain. Moreover, pure pixel-level mixing may lead to model confusion because of feature inconsistencies caused by distinct domain contexts.
In this paper, we explore UDA-SS for video and image scenarios within a unified framework to promote cross-pollination of ideas. As shown in Fig. 1(d), we propose a Quad-directional Mixup (QuadMix) method to address the above challenge for comprehensive domain gap bridging. To tackle complex video temporal feature shifts, we further propose a flow-guided saptio-temporal feature aggregation module for domain alignment, which can also be extended to images.
Firstly, QuadMix comprises comprehensive mixing paths in four directions designed for intra- and inter-domains. Unlike existing methods that struggle with insufficient inter-domain mixing caused by distinct point attributes, we smooth features within each domain to create more generalized intra-mixed domains, which are then merged to generate more enhanced inter-mixed domains. Technically, we generate adaptive category-aware video/image patch templates from two domains on the fly, which are then internally fused with their original domains to create intra-mixed source and target domains with generalized feature embeddings. Additionally, both pixel level and feature level patch templates are interactively fused with the intra-mixed domains to construct more enhanced inter-mixed source and target domains, as depicted in Fig. 7(e). Critically, we conduct the QuadMix at both spatial (temporal) pixel and feature levels to alleviate feature inconsistencies arising from diverse domain-wise contexts across quad-mixed domains. Consequently, it bridges domain gaps sufficiently within an enhanced feature embedding space, facilitating effective knowledge transfer from source to target domain.
Furthermore, for UDA-VSS, to better mitigate spatial and temporal gaps for enhanced pixel-level discrimination, we propose an adaptive spatio-temporal feature aggregation module guided by optical flow for domain alignment, which can also be extended to UDA-ISS. Typically, inherent feature gaps across source and target domains often lead to scattered distributions within each category and overlapping across categories, posing challenges to domain adaptation. To address this, we propose optical flow-guided spatial aggregation for precise fine-grained feature extraction, integrating temporal clues into spatial features to maintain temporal consistency. Following this, we conduct entropy-based temporal aggregation across consecutive frames to enhance overall temporal coherence. This module effectively utilizes temporal clues to extract and aggregate video features across domains, enabling fine-grained feature alignment for enhanced adaptation.
In a nutshell, our contributions can be summarized:
Through addressing the shared major challenge of underlying domain distribution shift, for the first time, we study UDA-SS for video and image from a unified view by comprehensive quad-mixing at pixel and feature levels.
To further tackle the complex spatio-temporal shift within videos, we propose a flow-guided spatio-temporal feature aggregation module for fine-grained alignment, which can also be extended to image scenarios.
Extensive experiments on four popular benchmarks show that we achieve superior results in unified scenarios. Particularly, we obtain mIoUs of 61.5, 56.2, 66.8, and 62.6 on Synthia-Seq Cityscapes-Seq, VIPER Cityscapes-Seq, GTAV Cityscapes, and Synthia Cityscapes, respectively. The former two are UDA-VSS benchmarks and the latter two are UDA-ISS benchmarks. Equipped with Transformer, we further improve mIoUs by 5.7, 2.0, 9.3, and 4.9, respectively, setting the new state-of-the-arts. Ablation studies and detailed analysis verify the effectiveness.
II Related Work
@@ II-A Unsupervised Domain Adaption for Image Segmentation
UDA-ISS has garnered significant attention in both theoretical and practical research. At the image level, the mainstream is to transfer image styles while preserving original semantics. Various approaches including cycle translation [25], patch adversarial [44], dual path learning [22], and text-to-image diffusion models [45] have been proposed. At the feature level, influential metrics and their variants [29, 30, 46] are developed to address distribution discrepancies across domains. Additionally, using prototypes to represent feature centroids has been crucial in pseudo-label denoising [47], divergence measurement [32], [48], and contrastive learning [23, 49]. At the output level, adversarial adaptation on low-dimensional logit space has been proposed in [50]. Other works, such as progressive confidence [51] and curriculum learning [52, 53], aim to refine target pseudo-labels within a structured output space [54]. Moreover, from the perspective of feature distribution, generating intermediate domains via one-way [19, 24, 23] or two-way pixel-level mixing [21, 22] to bridge feature gaps has yielded impressive results. However, distinct point attributes and feature inconsistencies are major challenges that often hinder effective knowledge transfer, which also exist in video scenarios. While UDA-ISS has been widely studied, UDA-VSS is still in its early stages. In this paper, we are the first to unify the study of these two tasks to address the shared major challenges. Notably, previous methods handle cross-domain features separately without interaction [19, 21, 22, 23, 24], whereas we enhance this by performing QuadMix at the feature level to alleviate feature inconsistencies.
II-B Video Semantic Segmentation
Video Semantic Segmentation (VSS) involves pixel-level classification for consecutive frames and has gained increasing attention powered by deep neural networks and densely labeled datasets [55, 56]. Mainstream works focus on leveraging temporal relationships to enhance segmentation accuracy. Techniques such as motion-guided feature warping [57] and prediction warping [58], along with temporal propagation via recurrent networks [59] or auxiliary networks [60], have been developed. Sun et al. [61] further propose a coarse-to-fine method to capture static and dynamic contextual features. To improve inference efficiency, Hu et al. [62] suggest using weight-sharing subnetworks across frames and reconstructing spatial features through attention mechanisms [63]. Recently, semi- and weakly-supervised VSS have gained increasing attention. Lao et al. [64] introduce a feature enhancement network that models both short-term and long-term temporal dependencies, while IFR [65] presents a feature reconstruction technique to reduce inner-video overfitting and utilizes partial labels to guide the learning of unlabeled frames.
However, existing VSS works heavily rely on dense labels, making it difficult to adapt to shifted and unlabeled target videos. In this paper, we study domain adaptation to address this issue. Notably, we leverage video temporal clues by (1) using optical flow to generate cross-frame consistent target pseudo-labels; (2) integrating temporal clues into spatial features for each domain; and (3) aggregating spatial features across consecutive frames for fine-grained domain alignment.
II-C Unsupervised Domain Adaptive Video Segmentation
@@
Research on video domain adaptation initially emerged for action recognition, emphasizing sparse classification of entire video clips [66, 67, 68, 69]. Recently, interest in UDA-VSS has surged due to the rising demand for dense video perception. Pioneering works [35, 6] employ adversarial learning with DCGAN [70] to extract domain-invariant video features. Guan et al. [6] further refine this by aligning uncertain target predictions with more confident adjacent ones. However, asynchronous issues with generator-discriminator convergence often cause training instability. Additionally, auxiliary clues involving geometric motion [36] or optical flow [37, 38] are introduced to generate pseudo-labels for self-supervised learning. However, they train the model on samples from each domain alternately but overlook domain interactions, leading to isolated learning that hampers adaptation to the target domain. Meanwhile, Wu et al. [71] explore transferring spatial knowledge from image to video domains while overlooking temporal feature gaps. Cho et al.[39] perform one-way input-level mixing from source to target but fail to address distinct point attributes across domains. In addition, they only align single-frame spatial features, heavily relying on a pre-trained model from [6] to generate initial pseudo-labels and initialize model parameters. In contrast, our method is trained end-to-end and further mitigates temporal shifts.
Based on the above analysis, we propose a unified method to tackle intra-domain distinct point attributes that hinder effective video/image domain gap bridging. For video scenarios, we further utilize temporal clues and propose flow-guided feature aggregation for fine-grained domain alignment. We are the first to explore UDA-VSS using advanced transformer networks.
@@ III-A Method Overview
UDA-SS transfers segmentation knowledge from a labeled source domain to a related but shifted target domain without pixel-level labels. In this paper, we unify the study of video and image scenarios to address shared challenges posed by the large feature gap, promoting a deeper understanding and efficient knowledge sharing. Naturally, the inputs differ between the two scenarios. For UDA-VSS, let with representing source adjacent frames alongside optical flow, and to be truth label of , where denotes a stack operation along time dimension. The target frames and optical flow are with but without label . For UDA-ISS, the source image and label are denoted as and while the target image is without label . The goal is to overcome the large domain gap and develop a model that effectively generalizes to the target domain.
In this paper, we propose a unified method named QuadMix to bridge and mitigate the significant video or image domain gap within an explicit feature space. As depicted in Fig. 2, to address the shared major challenge of fragmentary and insufficient gap bridging, QuadMix integrates four-directional paths designed for both intra- and inter-domains, operating at both input, feature, and label levels. Furthermore, we propose the adaptive spatio-temporal feature aggregation guided by optical flow for fine-grained domain alignment, which can also extend to image scenarios. To generate accurate pseudo-labels for self-supervised learning, for UDA-VSS, we input and from previous timestamp into the model to obtain , which is then warped using to generate temporally coherent , where denotes video temporal deviation. For UDA-ISS, the pseudo-labels are generated by a momentum network [72].
As analyzed earlier, current methods always struggle with distinct point attributes which result in insufficient bridging, whilst purely pixel-level mixing often leads to feature inconsistencies. To address this, as illustrated in Fig. 2(a), our QuadMix first adaptively generates category-aware video patch templates online from two domains (marked in yellow or blue). These templates are utilized to create two intra-mixed domains with smoother and generalized feature embeddings, addressing the feature compactness within each domain. Building on this, we perform interactive mixing to generate two more enhanced inter-mixed domains, serving as comprehensive intermediates for gap bridging. QuadMix not only involves spatial (frames and labels) and temporal (optical flow), but also spatio-temporal feature-level mixing. For UDA-ISS, QuadMix is similar while excluding temporal clues.
@@
Besides the spatial feature gap common to both scenarios, video data also faces a temporal gap. For UDA-VSS, we propose an adaptive flow-guided spatial-temporal feature aggregation module for fine-grained domain alignment (Fig. 2(b)). Spatial-level aggregation is guided by source/target pseudo masks adaptively online, while at the temporal level, features are further integrated based on logit entropy to enhance overall temporal coherence. Notably, each frame’s spatial feature is enriched with information from previous frames via optical flow propagation. Based on this, the aggregated features across domains are fine-grained aligned, thereby improving pixel-level discrimination for the target domain. This module can also be extended to UDA-ISS using spatial aggregation.

### v3

@@ Abstract
Unsupervised Domain Adaptive Semantic Segmentation (UDA-SS) aims to transfer the supervision from a labeled source domain to an unlabeled and shifted target domain. The majority of existing UDA-SS works typically consider images whilst recent attempts have extended further to tackle videos by modeling the temporal dimension. Although two lines of research share the major challenges – overcoming the underlying domain distribution shift, their studies are largely independent. It causes several issues: (1) The insights gained from each line of research remain fragmented, leading to a lack of holistic understanding of the problem and potential solutions. (2) Preventing the unification of methods and best practices across two scenarios (images and videos) will lead to redundant efforts and missed opportunities for cross-pollination of ideas. (3) Without a unified approach, the knowledge and advancements made in one scenario may not be effectively transferred to the other, leading to suboptimal performance and slower progress. Under this observation, we advocate unifying the study of UDA-SS across video and image scenarios, enabling a more comprehensive understanding, synergistic advancements, and efficient knowledge sharing. To that end, we explore the unified UDA-SS from a general domain augmentation perspective, serving as a unifying framework, enabling improved generalization, and potential for cross-pollination, ultimately contributing to the practical impact and overall progress. Specifically, we propose a Quad-directional Mixup (QuadMix) method, characterized by tackling intra-domain discontinuity, fragmented gap bridging, and feature inconsistencies through four-directional paths designed for intra- and inter-domain mixing within an explicit feature space. To deal with temporal shifts within videos, we incorporate optical flow-guided feature aggregation across spatial and temporal dimensions for fine-grained domain alignment, which is extendable to image scenarios. Extensive experiments show that our method outperforms the state-of-the-art works by large margins on four challenging UDA-SS benchmarks. Our source code and models will be released at https://github.com/ZHE-SAPI/UDASS.
@@
Unsupervised domain adaptation, semantic segmentation, unified adaptation, domain mixup.
I Introduction
Unsupervised Domain Adaptive Semantic Segmentation (UDA-SS) focuses on transferring pixel-level dense knowledge from a labeled source domain to an unlabeled target domain with distribution shift [1, 2, 3, 4]. The key challenge is all about how to tackle intrinsic distributional gaps between the two domains without target domain supervision. As typical, the literature is dominated by image UDA-SS with the aim to reduce the domain discrepancy through image alignment [5, 6, 7], metric minimization [7, 8, 9], self-supervised learning with pseudo-label [10, 11, 12], or regularization [13, 14]. This inspires the progress of video UDA-SS by additionally handling the temporal dimension via video alignment [15, 16], geometric-based cues [17, 18] for domain-invariant learning.
While the two UDA-SS tasks share the congenetic challenge, i.e., feature distributional gap across domains, their investigation is clearly separated. This could cause limitations like fragmented insights over tasks, lack of holistic understanding of the challenge, preventing the unification of techniques, and finally slowing the development progress.
Under the aforementioned insights, in this work, we advocate the unified investigation of image and video UDA-SS. To have a more generic solution, we explore the novel domain augmentation strategy by introducing a novel method, Quad-directional Mixup (QuadMix), capable of generating a diversity of intermediate domains within both source and target domains and across them within an explicit feature space, facilitated by an adaptive feature aggregation module for fine-grained domain alignment. Specifically, we generate adaptive category-aware sample patch templates from either (source or target) domain, followed by fusing them within the domain to create generalized intra-mixed domains. Concurrently, we further fuse those patches across the original domains interactively to construct more enhanced inter-mixed domains. In addition to common pixel-level mixing, feature-level mixing across the quad-mixed domains is also proposed to alleviate feature inconsistencies caused by domain-wise mixed-semantic discrepancies. Furthermore, to deal with persistent category-aware domain gaps, we design a unified feature aggregation module guided by spatial cues for image UDA-SS and spatio-temporal cues for video UDA-SS, enabling fine-grained domain alignment. Technically, our unified method goes beyond existing alternatives for image or video UDA-SS with limitations such as intra-domain discontinuity due to over distributed samples [10, 19, 16, 20], less generalizable feature distribution due to limited mixing directions [1, 2, 21, 5, 22], and feature inconsistencies due to diverse mixed-semantic contexts [22, 23].
We make the following contributions: (1) We for the first time investigate image and video UDA-SS from a unified perspective due to their shared major challenges. (2) We introduce a unifying UDA-SS framework, Quad-directional Mixup (QuadMix), which involves pixel- and feature-level mixing to facilitate comprehensive domain gap bridging and alleviate feature inconsistencies through four-directional paths for intra- and inter-domain mixing. For temporal shifts in videos, we incorporate optical flow-guided spatio-temporal feature aggregation, which is well-extensible to image scenarios. (3) As shown in Fig 1, extensive experiments on four image and video benchmarks show that our method sets the new state-of-the-art performance in comparison to prior works.
II Related Works
@@ II-A Unsupervised Domain Adaption for Image Segmentation
Image UDA-SS has garnered significant attention in both theoretical and practical research. At the image level, the mainstream is to transfer image styles while preserving original semantics. Various techniques including cycle translation [6], patch adversarial [24], dual path learning [22], and diffusion models [25] have been proposed. At the feature level, influential metrics and the variants [8, 9, 26] are developed to address distribution discrepancies across domains. Additionally, using prototypes to represent feature centroids is proposed for pseudo-label denoising [27], divergence measurement [12], [28], and contrastive learning [23, 29]. At the output level, adversarial adaptation on low-dimensional logit space has been proposed in [30]. Other works, such as progressive confidence [31] and curriculum learning [32, 33], aim to refine target pseudo-labels within a structured output space [34]. Moreover, from the perspective of feature distribution, generating intermediate samples [21, 1, 23, 5, 22] to mitigate domain gaps has yielded impressive results. However, the intra-domain discontinuity, less generalizable distribution, and feature inconsistencies are significant challenges that often hinder effective knowledge transfer to the target domain.
II-B Unsupervised Domain Adaptive Video Segmentation
Extending image segmentation, Video Semantic Segmentation (VSS) entails pixel-level discrimination for adjacent video frames [35, 36]. Mainstream works focus on enhancing segmentation accuracy, proposing methods involving feature warping [37, 38], recurrent networks [39], or auxiliary networks [40]. To improve inference efficiency, methods like attention-based spatial feature reconstruction [41] and weight-sharing subnetworks [42] have been proposed. However, they heavily rely on laborious dense labels, posing challenges in adapting to shifted and unlabeled target videos.
@@ II-B Unsupervised Domain Adaptive Video Segmentation
To address this, video UDA-SS has recently gained great interest. Pioneering works [15, 16] employ adversarial learning with DCGAN [43] to extract domain-invariant video features. Guan et al. [16] further improve this by aligning uncertain target predictions with more confident adjacent ones. However, asynchronous issues with generator-discriminator convergence often cause training instability. Additionally, auxiliary cues involving geometric motion [18] or optical flow [17, 19] are introduced to generate pseudo-labels. However, they train the model on samples from each domain alternately but overlook domain interactions, leading to isolated learning that hampers adaptation. Meanwhile, Wu et al. [44] explore transferring spatial knowledge from image to video domains while overlooking temporal feature gaps. Cho et al.[2] perform one-way pixel-level mixing from source to target but overlook the intra-domain discontinuity across domains. In addition, they obtain pseudo-labels offline using models in [16] and only align single-frame spatial features. In contrast, our method is trained end-to-end and holistically mitigates temporal shifts.
From a unified view, we consider the image to be a special case of the video without the temporal dimension. For video UDA-SS, we further leverage video temporal cues fully by (1) using optical flow to generate cross-frame consistent pseudo-labels; (2) integrating temporal cues to spatial features for each domain; and (3) aggregating video features of consecutive frames for fine-grained domain alignment. This is also the first work applying the vision transformer [45] to video UDA-SS.
II-C Intermediate Domains for Adaption
Current works achieved impressive results by addressing feature gaps with meticulously designed intermediate domains for images [21, 5, 22, 23, 1, 4] and videos [2]. However, the discontinuous distribution of features within each domain (intra-domain) always results in distinct point attributes [46], which makes current intermediates derived from such distinct domains with poorly generalized distributions [47]. Fig. 3 (a)-(d) presents a comparison of existing methods, distinguishing source features in green and target features in purple. Inspired by [48, 49], strategies like one-way mixing [21, 1, 23, 2] and two-way mixing [5, 22] extensively explore the integration of source patches into the target domain, or vice versa, aiming to bridge domain gaps. Nonetheless, t-SNE [50] feature visualization in Fig. 9 (a)(c)(d) reveals that the effect of existing intermediates across domains (inter-domain) is fragmented and insufficient, and persistent category-aware domain gaps continue to hinder effective knowledge generalization. Moreover, pure pixel-level mixing may lead to model confusion because of the feature inconsistencies caused by distinct domain-wise contexts.
@@ III Method
In this section, we start with an overview in Section III-A. Then, we elaborate on pixel-level mixing for intra- and inter-domains, as well as feature-level mixing across the overall quad-mixed domains in Section III-B to Section III-D. Next, Section III-E presents the flow-guided spatio-temporal feature aggregation module, and Section III-F covers the training loss.
@@ III-A Method Overview
Generally, current studies tend to address image and video UDA-SS tasks independently despite the shared challenges, leading to redundant efforts and missed opportunities for cross-pollination. In this paper, we tackle the more challenging task of unified UDA-SS. As depicted in Fig. 2, we propose a QuadMix method that integrates four-directional paths designed for both intra- and inter-domains, operating at spatial (frames and labels), temporal (optical flow of videos), and feature levels. Technically, the online generated category-aware video/image patch templates from two domains are internally fused with their original domains to create intra-mixed source and target domains with smoother and generalized feature embeddings. Additionally, both pixel- and feature-level patch templates are interactively fused with the intra-mixed domains to generate more enhanced inter-mixed source and target domains, serving as comprehensive and efficient intermediates for gap bridging (see Fig. 9 (e)). Our QuadMix is fairly practical and effective. A detailed comparison with previous works is shown in Fig. 3.
@@
Video is uniquely characterized by the temporal dimension as compared to image. To account for this intrinsic information, we propose an adaptive flow-guided spatial-temporal feature aggregation module for fine-grained domain alignment (Fig. 2 (b)). The spatial feature of each frame is enriched with temporal knowledge of previous frames via optical flow propagation. Based on this, spatial-level aggregation is guided by pseudo masks from the source and target domain adaptively. While at the temporal level, features are further integrated based on logit entropy to enhance overall temporal coherence. The aggregated features across domains are finely aligned in a fine-grained manner. It can be extended to image UDA-SS using spatial feature aggregation, as an image is essentially a special case of a video without the temporal dimension.
The main notations are listed as follows. For video UDA-SS, let with representing source frames alongside optical flow, and to be the truth label of , where denotes a stack operation along time dimension. The target frames and optical flow are and but without label . For image UDA-SS, the source image and label are denoted as and while the target image is without label . As for the pseudo-label generation, for the video scenario, and from previous timestamp are input into the model to obtain , which is then warped using to generate temporally consistent , where denotes the video temporal deviation. For the image scenario, pseudo-labels are generated by a momentum network [51].